OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

OpenStudy (anonymous):

2y-9

OpenStudy (anonymous):

-9+2y. Rearrange the first equation to get -x=9-2y, then times by -1 to get x=-9+2y

OpenStudy (anonymous):

or move x to other side to obtain 2y=9+x then move 9 to left side 2y-9=x if i see a - on one side i usally try to move it 1st